Mast Ch26 Multiple Choice Question 26.22 You completed Multiple Choice Question 26.22 Part A What is the frequency of an electromagnetic wave that has a wavelength of 300,000 km? less than 1 ㎐ O 1 Hz more than 1 Hz My Answers Give Up Correct
Answer #1

Given , λ = 300,000 km = 300,000*10^3 m = 3*10^8 m

Speed (v) = 3*10^8 m/s

Frequency (f) = v/λ = 3*10^8 /3*10^8 = 1 Hz

f = 1 Hz

Answer) option b)1 Hz